# Assessing Language Comprehension in Large Language Models Using Construction Grammar

### v1

@@ Appendix I Model Hyperparameters
We use the default hyperparameters for both GPT-4o and GPT-3.5, though we set the temperature of model responses to 0, to ensure maximal reproducibility of our work. All our experiments were done through the OpenAI API. The total cost of our experiments was approximately $250 USD.
Appendix J Llama Results
In addition to experiments with GPT-3.5 and GPT-4o we also ran additional experiments using LLaMA-3-8b-instruct and LLaMA-3-70b-instruct and have shared the results in the tables below. Overall, we observe that the trend of this dataset being difficult for LLMs holds.
The patterns we see Table 21 are similar to what we have seen with GPT:
Adding more in-context examples of construction improves performance.
Increasing the number of SNLI examples does not improve performance.
This reinforces our first conclusion that our dataset tests a different axis of semantic knowledge than more general datasets.
In Tables 2, 3 and 4 we see that the performance degrades compared to the performance on the first experiment, similar to the GPT models in the paper. This demonstrates that these tasks are also a difficult problem for LLaMA 3 and therefore act as illustrative examples of LLM capabilities.
In some cases we observe a slight difference in which Cxn/setting each model performs better in. We believe this is a reflection of their instruction fine-tuning data, which is likely to be different given that they have different creators. However, overall we observe that model performance improves with increasing parameter count and we see similar patterns in task difficulty for GPT and LLaMA models which are consistent with the findings in our paper.
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.59 | 0.74 |
| One-shot | CxNLI | 0.65 | 0.84 |
| Three-shot | CxNLI | 0.73 | 0.91 |
| One-shot | SNLI | 0.59 | 0.75 |
| Three-shot | SNLI | 0.58 | 0.74 |
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.59 | 0.74 |
| One-shot | CxNLI-challenge | 0.53 | 0.71 |
| One-shot | CxNLI-challenge, CxNLI | 0.66 | 0.84 |
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.38 | 0.52 |
| One-shot | CxNLI | 0.29 | 0.60 |
| Three-shot | CxNLI | 0.35 | 0.50 |
| One-shot | SNLI | 0.37 | 0.54 |
| Three-shot | SNLI | 0.39 | 0.57 |
| Cxn | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | |
| Comparative-Correlative | 0.27 | 0.5 |
| Let-Alone | 0.42 | 0.53 |
| Way-Manner | 0.45 | 0.44 |

### v2

@@ Appendix G Model Parameters and Hyperparameters
We use the default hyperparameters for all models including GPT-4o (unreported parameter size), GPT-3.5 (175B parameters), GPT-o1 (unreported parameter size), LLaMA 3 8B (8B parameters), LLaMA 3 70B (70B parameters) though we set the temperature of model responses to 0, to ensure maximal reproducibility of our work. Our GPT experiments were done through the OpenAI API. The total cost of our experiments was approximately $250 USD. Our LLaMA experiments were run through the replicate API and the experiments cost approximately $25 USD.
Appendix H Constructional Natural Language Inference Annotation Guidelines
Here we provide the exact instructions given to people to annotate the NLI datasets.
We have developed a dataset of sentences featuring different linguistic “constructions”—pairings of form and meaning. The constructions exemplified in this dataset range from purely substantive (the words filling the constructional slots are fixed), such as the Much-less construction, e.g., “He won’t eat shrimp, much less squid;” to purely schematic (the words filling the constructional slots can vary, but fulfill some general semantic and syntactic requirements), such as the Caused-Motion construction, e.g., “She blinked the snow off her eyelashes.” It’s okay if you aren’t familiar with this terminology or the idea of these constructions!
Task Overview
Your job is to read the sentences from this dataset, which are presented as the Premise in a set of triples for the Natural Language Inference (NLI) task. Also known as Recognizing Textual Entailment (RTE), NLI is the task of determining the inference relation between two (short, ordered) texts: entailment, contradiction, or neutral MacCartney and Manning (2008).
Premise: A man inspects the uniform of a figure in some East Asian country.
Hypothesis: The man is sleeping.
Then, you will fill in the Relation between the Premise and the Hypothesis, which indicates the kind of entailment between the two sentences. We are using numerical coding, listed below and in your annotation spreadsheet:
0 –
entailment – The hypothesis must be true given the premise.
1 –
neutral – The hypothesis may or may not be true given the premise.
2 –
contradiction – The hypothesis must not be true given the premise.
So for the example above, the correct answer would be:
2 – contradiction – If the man is inspecting a uniform, then it must not be true that the man is sleeping.
The two sentences describe the same scenario. Entities mentioned in both premise and hypothesis refer to the same thing; e.g., “the man” refers to the same individual. The hypothesis does not describe a different time.
If you encounter unfamiliar words, you may consult a dictionary. However, it is not expected or encouraged that you would have to “do research” into a topic in order to determine a relation between a premise and hypothesis. Instead, you should rely on common sense and your understanding of the words.
You will be completing these annotations in a spreadsheet like what is shown below, where there is a relation space available below a given premise/hypothesis pair. In the cell to the right of “relation”, you will provide the appropriate relation number. This space should include nothing except for the numbers 0, 1, or 2.
If you would like to note instances that are problematic, please add a Notes column to the right of Annotation Target and make your note as relevant to the right of the premise, hypothesis or relation.
There is also a space in column L where you can note the start and end times of each annotation session, or “sitting.” Please kindly track about how many judgments you are able to do in each sitting, so we can get a sense of how long the annotation task takes.
There are about 100 distinct NLI judgments or triples per annotation spreadsheet. Once you have completed all the relation annotations, please save and send the spreadsheet back to me: Claire.n.bonial.civ@army.mil.
Get your inference hat on. Happy annotating!